Here’s the last question I need help on. Thank you for anyone who solved this!!

If angle A and angle B are alternate interior angles solve for X given:
< A = 2x and B = x + 30


Answer by ikleyn(53619)

.

I solved this problem today 4 hour ago under this link

https://www.algebra.com/algebra/homework/Angles/Angles.faq.question.1179538.html

https://www.algebra.com/algebra/homework/Angles/Angles.faq.question.1179538.html


For your convenience, I copy-paste my solution here AGAIN.


--------------


Let me start from the proper definition


    +-------------------------------------------------------------------------+
    |   Alternate interior angles are angles formed when two parallel         |
    |   are intersected by a transversal.                                     | 
    |   The angles are positioned at the inner corners of the intersections   |
    |   and lie on opposite sides of the transversal.                         |
    +-------------------------------------------------------------------------+



One of the first theorem of Geometry states that alternate interior angles at two parallel lines are CONGRUENT, 
i.e. have equal angular measure.


It means that  A = B,   or


    2x = x + 30.


From this equation


    x = 30 degrees.


So, both angles A and B have equal measure of 2*30 = 60 degrees = 30+30 degrees.    ANSWER

Solved.

----------------

About parallel lines and associated angles see the lesson
    - Parallel lines
in this site.


///////////


Do not forget to post your  "THANKS"  to me for my teaching.